four times the larger of two consecutive integers is 94 more than three times the smaller... write the equation and solve?
Answer by ankor@dixie-net.com(22740)

four times the larger of two consecutive integers is 94 more than three times the smaller... write the equation and solve?
:
Let the two integers be: x, (x+1)
:
4(x+1) = 3x + 94
4x + 4 = 3x + 94
4x - 3x = 94 - 4
x = 90
:
:
Check this:
4(91) = 3(90) + 94
364 = 270 + 94